OpenStudy (anonymous):

Choose the equation below that represents the line that passes through the point (7, -2) and has a slope of -3. A. y - 2 = -3(x + 7) B. y + 7 = -3(x - 2) C. y + 2 = -3(x - 7) D. y - 7 = -3(x + 2)

OpenStudy (anonymous):

CAN SOMEONE HELP ME !!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

C. y + 2 = -3(x - 7)

OpenStudy (anonymous):

HOW DID YOU GET ?

OpenStudy (anonymous):

Through point slope form: \[(y-y _{1}) = m(x-x _{1}) \] The slope is -3: m = -3; and you pass through (7,-2); X1 = 7, Y1 = -2

OpenStudy (anonymous):

so plug in those values you get \[(y-(-2)) = -3(x-7)\]